A group of children was playing cards and misplace the five of hearts, if a card is selected at random from the remaining 51 cards. Find the probability that the card is a
a) club.
b) five.
c) five or club.
d) five of heart

Answer by Boreal(15235)

probability it is a club is 13/51
a 5 is 3/51 or 1/17
a five or a club is probability of both - conjoined probability (1/51)=(13+3-1)/51=15/51 or 5/17.
five of hearts is 0--it is missing.
five or a heart (if that is what was meant) is 3/51+(12/51), and there is no conjoined probability since the H5 is missing It is 15/51 or 5/17.